OpenStudy (anonymous):

What is the inverse of h? h(x) = 6x + 1

OpenStudy (anonymous):

substitute y for h(x) and rewrite y = 6x + 1. now swap x with y... so x = 6y+1 now solve for y

OpenStudy (anonymous):

can u do the rest?

OpenStudy (anonymous):

im stuck with x/6-1/6

OpenStudy (anonymous):

x = 6y + 1 -1 = 6y \[y =-\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

would the answer be.... 1/6(x-1)?

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[h ^{-1}(x)=-\frac{ 1 }{ 6 }x+1\]

OpenStudy (anonymous):

sorry was trying to get my kids to bed.